Compute the total cost per year of the following pair of expenses. Then complete the​ sentence: On an annual​ basis, the first set of expenses is​ _______% of the second set of expenses. Vern buys eight lottery tickets each week at a cost of ​$3 each and spends ​$850 per year on his textbooks.

On an annual​ basis, the money spent on lottery tickets is ______% of the money spent to buy textbooks.
Answer #1

Cost of 8 lottery ticket=3*8=24

Therefore cost in year,

ie, 24*52=$1248 on lottery tickets annually

And,$850 annually on textbooks.

1248 is 146.82%of850

the first set of expenses is​ ___146.82____% of the second set of expenses.

On an annual​ basis, the money spent on lottery tickets is __146.82____% of the money spent to buy textbooks.